A hot dog vendor faces a daily demand curve of Q=1800-15P.

If the vendor has been selling 300 hot dogs each day. How much revenue has he been collecting?


1
Expert's answer
2021-05-18T19:35:42-0400

Given Q=300, the price P per hot dog is;

Q=1800−15P

300=1800−15P

15P=1500

P=100

Revenue=P×Q=100×300=30,000


The vendor is collecting revenue equal to $30,000.